OpenStudy (mikeymaximum):

Yolanda's recipe called for 4 pints of chicken stock. How many cups of chicken stock did Yolanda need? (1 pint = 2 cups) A. 2 cups B. 4 cups C. 6 cups D. 8 cups

OpenStudy (bannicularf):

d

OpenStudy (a1234):

1/2 = 4/x x = 8

jimthompson5910 (jim_thompson5910):

1 pint = 2 cups ... given 4*(1 pint) = 4*(2 cups) ... multiply both sides by 4 since we have 4 pints (4*1) pints = (4*2) cups 4 pints = 8 cups